Calculate the double integral ||(x + 3 y) dA where R is bounded by y = Vx and y = x
Answer #1

+ 3) A, R: 4 , 4 = x- 1 =22 -yA /C ) (0) I d - Stop - 2 ] 8+I5-S-6 - + - - 중 = SIS-5-6 SAIN